OpenStudy (anonymous):

slove for x 5(3x=4)-6x+3(3x+2)

OpenStudy (anonymous):

Double check the equation to see if you wrote it down correctly, please.

OpenStudy (anonymous):

i wrote it down right

OpenStudy (anonymous):

15x=20-6x+9x+6 15x+6x=26 21x=26 x=26/21

OpenStudy (anonymous):

thankz

OpenStudy (anonymous):

You're welcome!